OpenStudy (anonymous):

Given that a/3 and b/3 are the roots of the quadratic equation mx(2x-3)=10x-n,calculate the values of m and n if a+b=19/2 and ab=45/2. @rational

OpenStudy (anonymous):

\[2mx^2-3mx-10x+n=0\]\[2mx^2-(3m+10)x+n=0\]

OpenStudy (rational):

Yes use the vieta's formulas sum of roots : \[\frac{a}{3}+\frac{b}{3}=3m+10\] product of roots: \[\frac{a}{3}\cdot\frac{b}{3}=n\]

OpenStudy (anonymous):

\[\frac{ a+b }{ 3 }=3m+10\]\[\frac{ ab }{ 9 }=n\]

OpenStudy (anonymous):

sholud it be \[SOR=-\frac{ b }{ a }\]\[POR=\frac{ c }{ a }\]? @rational

OpenStudy (rational):

Oops yes, my mistake!

OpenStudy (rational):

thnks for catching :)

OpenStudy (rational):

So after fixing that mistake, do we get sum of roots : \[\frac{a}{3}+\frac{b}{3}=\frac{3m+10}{2m}\] product of roots: \[\frac{a}{3}\cdot\frac{b}{3}=\frac{n}{2m}\]

OpenStudy (anonymous):

\[\frac{ a+b }{ 3 }=\frac{ 3m+10 }{ 2m }\]\[\frac{ ab }{ 9 }=\frac{ n }{ 2m }\]

OpenStudy (anonymous):

should we substitute a+b=19/2 and ab=45/2 into the equations? @rational

OpenStudy (rational):

Yes

OpenStudy (anonymous):

i got \[m=-\frac{ 15 }{ 4 } and~n=-\frac{ 75 }{ 4 }\]but the answer from the book say it is \[m=3,n=15\]

OpenStudy (radar):

Check your work, you will get 20m = 60 or m=3

OpenStudy (anonymous):

\[-\frac{ 3m+10 }{ 2m }=\frac{ 19 }{ 6 }\]\[-18m-60=38m\]\[m=-\frac{ 15 }{ 4 }\]

OpenStudy (radar):